OpenStudy (anonymous):

Write the explicit formula that represents the geometric sequence -2, 8, -32, 128

OpenStudy (misty1212):

HI!!

OpenStudy (anonymous):

hi

OpenStudy (misty1212):

what do you multiply \(-2\) by to get \(8\)? this is another way of asking what is \(8\div(-2)\)

OpenStudy (anonymous):

-16

OpenStudy (misty1212):

hmm no

OpenStudy (anonymous):

-9

OpenStudy (priyar):

no..

OpenStudy (anonymous):

-4

OpenStudy (priyar):

what is 8/2 ?

OpenStudy (anonymous):

4

OpenStudy (priyar):

yeah! so here it just has a minus sign extra coz it is..-8/2 so it will be -4

OpenStudy (priyar):

ok?

OpenStudy (anonymous):

yeah

OpenStudy (priyar):

so r =-4

OpenStudy (priyar):

are u familiar with this notation?

OpenStudy (anonymous):

no i am not

OpenStudy (priyar):

r stand for common ratio of a G.P..

OpenStudy (priyar):

the sequence goes like: a,ar,ar^2,ar^3 and so on..got it?

OpenStudy (dayakar):

let a1= -2 a2= 8 a3 = -32 a4 = 128 find ratio a2/a1 = 8/-2 = a3/a2 = a4/a3 = if u get all above ratios are equal the above sequence is in geometric sequence

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

im still lost . do i just do the 4/3= is that what i have to do.

OpenStudy (dayakar):

no a1 means first term

OpenStudy (dayakar):

a2/a1 = 8/-2 = -4 a3/a4 = -32/8 =

OpenStudy (dayakar):

is it clear

OpenStudy (anonymous):

i think so. thank you

OpenStudy (dayakar):

here u get ratio = -4 this is called common ratio denoted with 'r'

OpenStudy (anonymous):

yeah they all come out as -4

OpenStudy (dayakar):

general form of the sequence is 

OpenStudy (anonymous):

okay thank i will get it

OpenStudy (dayakar):

:)

OpenStudy (anonymous):

:)